OpenStudy (anonymous):

help meeeee again, please? Let f(x) = (x2 - 2x)2 - 23(x2 - 2x). Find all x such that f(x) = -120. a. 4,5 b. 8,15 c. -2,-3,8,15,4,5 d. -2,-3,4,5

jhonyy9 (jhonyy9):

a. will be (16-8)2 -23(16-8)=64-184 = -120 so is right

jhonyy9 (jhonyy9):

225-23*15=225-345=-120 is right

jhonyy9 (jhonyy9):

so a. is right all